3b) A bank pays you 4.5% interest per year. In addition, you receive |100 as bonus at

the end of the year (after the interest is paid). Find a recurrence for the amount of

money after n years if you invest |2000.


1
Expert's answer
2021-04-15T07:33:43-0400

Let r=0.045 is an interest rate, S0=2000 the initial amount of money, Sn - the amount of money after n years, and b=100 is the annual bonus. Then we have a recurrent equation:

"S_n=(1+r)S_{n-1}+b"

"S_n-S_{n-1}=r(S_{n-1}+b\/r)"

"(S_{n}+b\/r)-(S_{n-1}+b\/r)=r(S_{n-1}+b\/r)"

"S_{n}+b\/r=(1+r)(S_{n-1}+b\/r)"

The last equation shows that the sequence "S_{n}+b\/r" is a geometric progression, and we can write

"S_{n}+b\/r=(1+r)^n(S_{0}+b\/r)"

"S_{n}=(1+r)^n(S_{0}+b\/r)-b\/r"

"S_{n}=(1+0.045)^n(2000+100/0.045)-100/0.045=4222.22\cdot1.045^n-2222.22"

Answer. "S_{n}=4222.22\cdot1.045^n-2222.22"